# Interpretable Spatio-Temporal Features Extraction based Industrial Process Modeling and Monitoring by Soft Sensor

### v1

@@ IV-D2 The importance of far topological alignment
By replacing the far topological alignment layer with a CNN model, we emphasize the necessity of the far topological alignment layer in STCIM. Since power generation in a power plant is a very complex industrial process, the input variables are coupled with each other. At the same time, the inertia of each device causes the spatial and temporal data to be misaligned, affecting the accuracy of modeling. Table VI describes how the evaluation metrics change when the far topological alignment is replaced. All evaluation metrics become worse at both test 1 and test 2, especially in the $y$ prediction. The lack of far topological alignment makes the model focus more on local features, while ignoring the potential dependencies of far topological correlations of variables, reducing the accuracy of modeling and monitoring, which is fatal for a large and complex industrial process.
TABLE VI: Evaluation metrics of experiments in a real-life dataset without far topological alignment
[…]
@@
Test 1 0.97 0.42 0.06 0.90 0.51 0.32 0.46 1.38 1.45
Test 2 1.78 0.13 0.09 1.08 0.32 0.09 2.93 1.56 0.82
IV-D3 Sensitivity Analysis of hyper-parameters $\gamma$
The impact of hyper-parameters $\gamma_{1}$ and $\gamma_{2}$ on performance of STCIM are tested in this part. Figure 8 shows the performance of STCIM using different $\gamma$. When the state equations are not considered in the loss function, the output $y$ and state variables are unable to align with actual data. Additionally, the utilized part of the state equations can partly guarantee the fit of state variables. The complexity of the expression determines which state variable is automatically fitted by the neural network. When $\gamma_{1}=1$ and $\gamma_{2}=1$, there is more oscillation in Figure 8d compared with Figure 6b since more loss about state variables are considered in physics-informed interpretable latent space block.
(a) Test 2 for the real-life dataset with $\gamma_{1}=0$ and $\gamma_{2}=0$
(b) Test 2 for the real-life dataset with $\gamma_{1}=0$ and $\gamma_{2}=1$
(c) Test 2 for the real-life dataset with $\gamma_{1}=1$ and $\gamma_{2}=0$
(d) Test 2 for the real-life dataset with $\gamma_{1}=1$ and $\gamma_{2}=1$
Figure 8: Sensitivity analysis of hyper-parameters $\gamma_{1}$ and $\gamma_{2}$
@@ V Conclusion
As modern industrial processes become complex, general soft sensor methods are gradually unable to meet the requirements. This paper introduces a novel soft sensor method termed spatio-temporal consistent and interpretable model (STCIM). We designed a far topological spatio-temporal consistency extraction block to capture and align temporal and spatial features from raw data. Subsequently, by giving physical meanings to latent variables and integrating them in a physics-informed discrete state-space loss function, the physics-informed interpretable latent space block maps the spatio-temporal features to an interpretable state-space to improve the feature interpretability. The primary outcome underscores the STCIM’s capability to effectively extract and align the spatio-temporal features and the far topological structure. The analytical experiments underscore that STCIM has robust generalization performance across diverse operational conditions, particularly when the training dataset exhibits significant fluctuations in both outputs and state variables, thereby enabling informed and reliable decision-making among power domain experts. The ablation experiments highlight the critical significance of adaptive position encoding, the far topological alignment, and the hyper-parameters of the loss function.

### v2

@@ IV-D2 The importance of far topological alignment
Table V describes how the evaluation metrics change when the far topological alignment is replaced by a CNN model. Different from the comparisons in Table IV, the performance of STCIM at both Test 1 experiment is greatly affected, especially in the $y$ prediction and $z_{1}$ estimation, which means that for complex thermal power plants and coupled variables, global and topological features are more important. In Test 2 experiment, the performance becomes worse, due to the change of operating point. The lack of far topological alignment makes the model focus more on local features, while ignoring the potential dependencies of far topological correlations of variables, reducing the accuracy of modeling and monitoring, which is fatal for a large and complex industrial process.
TABLE V: Evaluation metrics of experiments in a real-life dataset without far topological alignment
Datasets MSE MAE MAPE
@@
Test 1 0.97 0.42 0.06 0.90 0.51 0.32 0.46% 1.38% 1.45%
Test 2 1.78 0.13 0.09 1.08 0.32 0.09 2.93% 1.56% 0.82%
IV-D3 Sensitivity analysis of hyper-parameters $\gamma$ in Test 2
The impact of hyper-parameters $\gamma_{1}$ and $\gamma_{2}$ on performance of STCIM are tested. Fig. 6 shows the performance of STCIM using different $\gamma$. When the discrete state-space loss (Fig.(6a)) is not considered in final loss function, STCIM becomes a supervised output and state variables forecasting like VCACNN and the forecasting is unable to align with actual data even though the trend is similar. In Figs.(6b) and (6c), the utilized part of the discrete state-space loss can partly guarantee the fit of state variables. Due to the strong correlation between the state variables and the output, actual fuel and pressure of steam follow the trend of output power, which is different from the actual values. In Fig.(6d), the pressure of steam follows the trend of the real values, due to all the discrete state-space equations are used as part of the loss function.
Figure 6: Sensitivity analysis of hyper-parameters $\gamma_{1}$ and $\gamma_{2}$ in Test 2 using thermal power plant dataset. (a) $\gamma_{1}=0$ and $\gamma_{2}=0$. (b) $\gamma_{1}=1$ and $\gamma_{2}=0$. (c) $\gamma_{1}=0$ and $\gamma_{2}=1$. (d) $\gamma_{1}=1$ and $\gamma_{2}=1$.
TABLE VI: Evaluation metrics of experiments with different input dimension
Datasets MSE MAE MAPE
Type y{y} $z_{1}$} $z_{2}$} y{y} $z_{1}$} $z_{2}$} y{y} $z_{1}$} $z_{2}$}
H=5 0.82 0.06 0.11 0.62 0.22 0.13 0.28% 0.47% 0.58%
H=10 0.56 0.03 0.09 0.51 0.14 0.07 0.17% 0.41% 0.40%
H=20 0.49 0.02 0.03 0.46 0.11 0.05 0.14% 0.35% 0.33%
IV-D4 The impact of input dimension $H$ in Test 2
Table VI discusses the impact of input dimension $H$ on the performance of STCIM ($Input\in R^{H\times W\times C}$). Due to the limited computational resources, we test three different $H$. When the backtracking is small, the advantages of STCIM is hard to see. The inner far topological consistent spatio-temporal feature is replaced by the local spatio-temporal feature, making the performance of STCIM comparable to that of CNN-based models. As $H$ increases, the performance of STCIM is further enhanced. However, considering the computational complexity of STCIM in attention mechanism, a larger $H$ will bring more burden to the device.
@@ V Conclusion
As modern industrial processes become complex, general soft sensor methods are gradually unable to meet the requirements. This paper introduces a novel soft sensor method termed spatio-temporal consistent and interpretable model (STCIM). We designed a spatio-temporal consistent feature extraction block to capture and align local and global temporal and spatial features from raw data. Subsequently, by giving physical meanings to latent variables and integrating them in a physics-informed interpretable latent space loss function, the latent space maps the spatio-temporal consistent features to an interpretable latent variables to improve the model interpretability. The primary outcome underscores the capability to effectively extract and align the spatio-temporal features and the interpretability of STCIM. The analytical and comparative experiments underscore that STCIM has robust generalization performance across diverse operational conditions, particularly when the training dataset exhibits significant fluctuations in both outputs and state variables, thereby enabling informed and reliable decision-making among power domain experts. The ablation experiments highlight the critical significance of adaptive position encoding, the far topological alignment, and the hyper-parameters of the loss function.